OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <6, 4>, v = <-9, 8>

OpenStudy (dumbcow):

use dot product \[u*v = (6*-9)+(4*8) = -22\] vectors are orthogonal only if dot product is zero

OpenStudy (anonymous):

so its not orthogonal,, its parallel or neither then right?

OpenStudy (dumbcow):

vectors are only parallel if they are proportional ..... <1,2> and <3,6> so answer is neither

OpenStudy (anonymous):

ohh ! ok thank you, do u have time to help me with one more?

OpenStudy (dumbcow):

ok

OpenStudy (anonymous):

Given that P = (-5, 5) and Q = (-13, 6), find the component form and magnitude of 2PQ

OpenStudy (anonymous):

so far i think its, <-16, 2> now i just have to find the magnitude

OpenStudy (dumbcow):

hmm you can't do the cross-product of 2-d vectors and get a vector ?

OpenStudy (anonymous):

huh?

OpenStudy (dumbcow):

what do they mean by PQ?

OpenStudy (anonymous):

im not really sure thats why im comfused

OpenStudy (anonymous):

im stuck on this problem and this one Find the fifth roots of 243(cos 240° + i sin 240°).

OpenStudy (dumbcow):

to find 5th root .... take 5th root of 243 and divide angle by 5

OpenStudy (anonymous):

i have to find the five roots to that equation i think?

OpenStudy (dumbcow):

right \[= 3(\cos (48+k \pi)+i \sin(48 +k \pi))\] where k = 0,1,2,3,4

OpenStudy (anonymous):

yeah, so what else would i do then from there?

OpenStudy (dumbcow):

sorry , pi should be 180 since you are in degrees oh wait i am wrong...forgot something \[= 3(\cos (\frac{240+180k}{5})+i \sin(\frac{240+180k}{5}))\]

OpenStudy (anonymous):

so thats the answer? now im comfused? what would i show as steps?

OpenStudy (dumbcow):

plug in each value for k and those are the 5 roots

OpenStudy (anonymous):

oh so plug the 0,1,2,3,4 into where the k is and thats it?

OpenStudy (dumbcow):

yeah

OpenStudy (anonymous):

how should i plug in the isin in the calculator

OpenStudy (anonymous):

i just get decimals when i plug them in?